answersLogoWhite

0

The probability of rolling a 7 on a pair of 2 dices is: P(7) = 6/36.

The probability of rolling a 11 on a pair of 2 dices is: P(11) = 2/36.

The probability of rolling 7 or 11 is the sum of both single probabilities:

P(7 or 11) = 6/36 + 2/36 = 8/36 = 2/9 = 0.2222... ≈ 22.2%.

Explanation.

When rolling a pair of fair dice there are 36 possible equiprobable events or

outcomes. They are:

(1,1), (1,2), (1,3), (1,4), (1,5), (1,6),

(2,1), (2,2), (2,3), (2,4), (2,5), (2,6),

(3,1), . . . . . . . . . . . . . . . . . , (3,6),

(4,1), . . . . . . . . . . . . . . . . . , (4,6),

(5,1), . . . . . . . . . . . . . . . . . , (5,6),

(6,1), . . . . . . . . . . . . . . . . . , (6,6).

Each of these 36 events has the probability of happening of 1/36.

For the case of the sum of both dice to be 7, there are 6 possible outcomes, so the

probability of the event sum equals 7 is: P(7) = 6/36 = 1/6.

For the case of the sum of both dice to be 11, there are 2 possible outcomes, (6,5)

and (5,6), so the probability of the event sum equals 11 is: P(11) = 2/36 = 1/18.

User Avatar

Wiki User

∙ 14y ago

What else can I help you with?

Related Questions

What is the probability of rolling a sum of 11 with 2 dice?

The probability of rolling a sum of 11 with 2 dice is: P(11) = 1/18.For explanation see answer to question: "What is the probability of rolling 7 or 11 with 2 dice?".Read more: What_is_the_probability_of_rolling_7_or_11_with_2_dice


What is the probability of rolling a 7 or 11 with two dice?

The probability of rolling a 7 with 2 dice is 6/36; probability of rolling an 11 is 2/36. Add the two together to find probability of rolling a 7 or 11 which is 8/36 or 2/9.


What is the probabilities of rolling a 2 with two dice?

The probability of rolling a sum of 2 is 1/36 The probability of rolling the value 2 on one die or the other (or both) is 11/36


When two dice are rolled what is the probability of not getting a sum of 12?

The probability of rolling 12 with 2 dice is 1 in 36. The probability of not rolling 12 with 2 dice is 35 in 36.


If two fair dice were rolled what is the probability of them rolling a sum of 11?

x123456123456723456783456789456789105678910116789101112Therfore, there is a 2 in 36 (or 1 in 18) chance of rolling a total of 11.


What is the probability of having the same number when rolling a dice twice?

If it is a regular dice then the probability is 3/6 that is 1/2


What is the probability of rolling the sum of 10 with 2 dice?

Well there is 36 different possibilities with rolling 2 6 sided dice. The probability of rolling the sum of 10 with 2 die is 4/36 or 1/8 chance.


What is the probability of rolling an odd number with a dice?

The probability for a normal die is 1/2.


What is the probablity of rolling the sum of 12 on 2 dice?

There are 36 outcomes for rolling 2 dice, and there is 1 way that a 12 can occur which is 6,6. So, the probability of rolling the sum of 12 on 2 dice is 1/36.


What is the probability of rolling a even number on 1 dice?

1/2.


What is the probability of rolling the sum of 9 with 2 dice?

1 in 9


What is the probability of rolling two numbers whose sum is 2 equal?

For 6 sided dice, there is only 1 way to get a 2: (1,1). There are 36 outcomes rolling 2 dice; so the probability of rolling two numbers whose sum is 2 is 1/36.